Find all exact solutions of: tanxsinx - sinx = 0
Answer by jim_thompson5910(35256)

tan(x)sin(x)-sin(x)=0 ... Start with the given equation

sin(x)(tan(x)-1)=0 ... Factor out the GCF sin(x)

sin(x)=0 or tan(x)-1=0 ... Set each factor equal to zero

Solve the first equation:

sin(x)=0 ---> x = 0 + 2pi*n ... or ... x = pi + 2pi*n


Solve the second equation:

tan(x)-1=0=0 ---> tan(x)=1 ----> x = pi/4 + pi*n

So the solutions are

x = 0 + 2pi*n ... or ... x = pi + 2pi*n ... or ... x = pi/4 + pi*n

where "n" is an integer.